Daphne Company reported the following results from last year's operations:

Sales

$13,200,000

Variable expenses

8,490,000

Contribution margin

4,710,000

Fixed expenses

3,654,000

Net operating income

$1,056,000

Average operating assets

$6,000,000

At the beginning of this year, the company has a $1,000,000 investment opportunity with the following characteristics:

Sales

Contribution margin ratio

$3,400,000

50% of sales

Fixed expenses

$1,496,000

The company's minimum required rate of return is 17%.

Required:

    1. What was last year's residual income?
    1. If the company pursues the investment opportunity and otherwise performs the same as last year, what will be the overall residual income this year? (provide calculations showing how you arrived at avg operating assets and residual income for new investment)
    2. If Jersey Company's CEO earns a bonus only if residual income for this year exceeds residual income for last year, would the CEO pursue the investment opportunity?
Answer #1

Residual income = Net operating income-Minimum Operating income

a) Residual income = 1056000-(6000000*17%) = 36000

b) Residual income = (1056000+204000)-(7000000*17%) = $70000

c) Yes, Company should pursue this investment opportunity